OpenStudy (anonymous):

HELP PLEASE! METAL. Find all solutions to the equation. sin2x + sin x = 0

OpenStudy (anonymous):

sin 2x + sin x = 0 2.sin x.cos x + sin x = 0 sin x(2 cos x + 1) = 0 2 cos x + 1 = 0 cos x = -1/2 x = 120 degrees

OpenStudy (anonymous):

How did you get that Answer?

Directrix (directrix):

Is the solution set supposed to come from the set of all real numbers? Did the instructions mention that?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

when no, but I assume so.

Directrix (directrix):

At this point in the work above, the Zero Product Property should have been used. sin x(2 cos x + 1) = 0 sinx x = 0 Or cos x + 1 = 0 sin x = 0 Or cos x = -1 @lalithavasanth

OpenStudy (anonymous):

That's it?

OpenStudy (anonymous):

but cos x should be equal to -1/2...?

OpenStudy (anonymous):

Is that it?

Directrix (directrix):

@clairvoyant1 What does this mean in your text? >>Find all solutions to the equation. We need to know over what domain we are solving the equation.

OpenStudy (anonymous):

sin squared of x + sin x = 0 ?

Directrix (directrix):

And, we need to know if by sin2x + sin x = 0 , you meant to convey: